# Preferences Evolve And So Should Your Bandits: Bandits with Evolving States for Online Platforms

### v2

@@ B.3 Generalization for Unknown Replenishing Arm
Set $\varepsilon,\delta,M$ as stated in Theorem B.3. Initialize rounds $t=1$. /* Explore IV rewards and build their estimators: $\{\widehat{r}_{i}\}_{i\in[K]}$/
@@
2       Initialize reward estimate $\widehat{r}_{i}=0$. for blocks $j\in[M]$ do // Restore the state to at least $b_{z}-\varepsilon$
             Choose an arm $z\in[K]$ uniformly at random.   // $z$ = benchmark arm for state.
@@
4                   Play arm $z$. Update $t\leftarrow t+1$.
            Play arm $i$, observe reward $R_{j}^{i}$, and update: $\widehat{r}_{i}\leftarrow\widehat{r}_{i}+\frac{R_{j}^{i}}{M}$.   // Play $i$ when $q\approx b_{z}-\varepsilon$.
@@
7       Initialize state estimator $\widehat{v}_{i}=0$. for pulls $1,\dots,N(\lambda)$ do
8             Play arm $i$. Update $t\leftarrow t+1$.
@@
             Play arm $i$, observe reward $S_{j}^{i}$, and update: $\widehat{v}_{i}\leftarrow\widehat{v}_{i}+\frac{S_{j}^{i}}{M}$.   // Play $i$ when $q\approx b_{i}$
10
@@
Play arm $i_{R}$ for $N_{R}$ rounds, updating $t\leftarrow t+1$ after each one.  // Restore state to at least $1-\varepsilon$
@@ Appendix C Appendix for Section 4
Input. $\eta=0.95\sqrt{\frac{\log K}{KT}}$, $\gamma=1.05\sqrt{\frac{K\log K}{T}},\beta=\sqrt{\frac{\log(K\delta^{-1})}{KT}},\forall\delta\in(0,1)$ Initially at time $t=1$, let $p_{1}$ be the uniform distribution over $[K]$. for $t\in[T]$ do
       Choose an arm $I_{t}\in[K]$ from probability distribution $p_{t}$. Observe reward $g_{I_{t},t}\sim\textsf{Bern}(r_{I_{t}}q_{t})$. For each arm $i\in[K]$ compute the estimated biased gain:
@@
and update the estimated cumulative gain: $\tilde{G}_{i,t}=\sum_{s\in[t]}\tilde{g}_{i,s}.$ Compute the new probability distribution over arms $p_{i,t+1}$:
@@ E.1 Noise-Perturbed Model
Note, that even in the noise-perturbed model the $q_{t}$ part of the state is still defined deterministically and Lemma 3.5 (and Corollary 3.6) still hold. Thus, after playing an arm $i$ repeatedly for $N(\lambda)$ rounds we get $|q_{t}-b_{i}|\leq\varepsilon$. Algorithm 2 plays replenishing arm $i_{R}$ for $N(\lambda)$ rounds and so $q_{t}\geq 1-\varepsilon$. After the state has converged to $b_{i_{R}}$, the algorithm starts building the estimators. While noise does not affect the $q_{t}$ part of the state, it does affect $\tilde{q}_{t}$ which in turn is where the estimators are built from. More specifically, for the $\hat{r}_{i}$ estimators, Algorithm 2 after playing $i_{R}$ for $N(\lambda)$ rounds, takes a sample of arm $i$; hence, the sample came from $\textsf{Bern}(r_{i}\cdot(1-\varepsilon+\nu_{t}))$. As a result, Lemmas 3.7 and 3.9 change as follows:
@@
The intuition and analysis of this case bears similarities with the analysis for the case where $\lambda=1-\epsilon$ (Section 5).
If the noise added to a round $\nu_{t}$ is “small enough” (specifically, if $\nu_{t}\leq\sigma\sqrt{\log(T/\delta)}$ for $\sigma<\sqrt{T}$), then the analysis is identical to the case where $\lambda=1-\epsilon$. If the noise added to a round $\nu_{t}$ is greater than $\sigma\sqrt{\log(T/\delta)}$, then in the worst case the expected reward can be affected by a $\sigma T$ factor in total. Putting everything together (and using the Hoeffding bound of Equation 27) we get the result. ∎
@@ E.2 Unknown $\lambda$
In this section, we present the analysis when $\lambda$ is unknown. We first present the Unknown General $\lambda$ which will be used in the main algorithm of the unknown $\lambda$. In Algorithm 7 we get as an input a $\tilde{N}(\lambda)$ and two arms $i,j$. Then, we use a technique from Algorithm 2 to learn $r_{i}b_{i}$ and $r_{i}q_{ALT_{i,j}}$, which will defined below. After that, we solve an equation and if $\tilde{N}(\lambda)$ is near the real $N(\lambda)$ we can get a good estimator $\hat{\lambda}$ for $\lambda$.
Input. $\tilde{N}(\lambda)$, arm $i$, arm $j$. Choose $\varepsilon$, $\delta$, $M$ based as in $\tilde{N}(\lambda)$ and Theorem 3.1 /* Build estimators for $r_{i}b_{j}$/
@@
2       for pulls $\in[\tilde{N}(\lambda)]$ do
             Play arm $j$.   // $q_{t}$ near $b_{j}$
@@
      Play arm $i$, observe reward $R_{i,b_{j}}$ and update: $\hat{r}_{i,b_{j}}=\hat{r}_{i,b_{j}}+R_{i,b_{j}}/M$.   // take a sample of $r_{i}\cdot b_{j}$
@@
5 for blocks $\in[M]$ do
6       for pulls $\in[\tilde{N}(\lambda)]$ do
             Play arm $j$.   // $q_{t}$ near $b_{j}$
7
8      for pulls $\in[\tilde{N}(\lambda)]$ do
             Play arm $i$.   // $q_{t}$ near $b_{i}$
9
      Play arm $i$, observe reward $R_{i,i}$ and update: $\hat{r}_{i,i}=\hat{r}_{i,i}+R_{i,i}/M$.   // take a sample of $r_{i}\cdot b_{i}$
10
/* Build estimators for $\frac{r_{i}b_{j}-(1-\lambda)r_{i}b_{i}}{2-\lambda}$/
11 for blocks $\in[M]$ do
12       for pulls $\in[\tilde{N}(\lambda)]$ do
@@ 13             Play arm $i$
14      for pulls $\in[\tilde{N}(\lambda)]$ do
15             Play arm $i$ Play arm $j$
      Play arm $i$, observe reward $R_{i,j}$ and update: $\hat{r}_{i,j}=\hat{r}_{i,j}+R_{i,j}/M$.   // take a sample of $r_{i}\cdot\frac{b_{j}+(1-\lambda)}{2-\lambda}$
16
Solve $\hat{\lambda}=\frac{\hat{r}_{i,i}+\hat{r}_{i,j}-2\hat{r}_{i,j}}{\hat{r}_{i,i}-\hat{r}_{i,j}}$ Output. $\hat{\lambda}$.
ALGORITHM 7 B-DES Unknown General $\lambda$
Formally, in this section we will be proving the following.
Theorem E.2.
Algorithm 8 achieves regret:
$R_{\textsc{DES}}(T)=\begin{cases}\widetilde{\mathcal{O}}\left(K^{1/3}T^{2/3}\right)&\textrm{for }\lambda\in[0,\Theta(1/T^{2})]\\ \mathcal{O}\left(T^{b/a}\right)&\textrm{for }\lambda=T^{-a/b}\textrm{ and }\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}>\frac{\sqrt{\log T}K^{1/3}}{T^{1/3}}\\ (1-1/e)\textsf{OPT}&\textrm{for }\lambda=\Theta(1/T).\textrm{ and }\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}>\frac{\sqrt{\log T}K^{1/3}}{T^{1/3}}\\ \widetilde{\mathcal{O}}\left(K^{1/3}T^{2/3}\right)&\textrm{for }\lambda\in(\Theta(1/T^{2}),1]\textrm{ and }\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}\leq\frac{\sqrt{\log T}K^{1/3}}{T^{1/3}}\\ \mathcal{O}\left(\left(\frac{K\log(T)\log(\lambda)}{\log(1-\lambda)}\right)^{1/3}\cdot T^{2/3}\right)&\textrm{for }\lambda\in(O(K/T)^{1/3},1]\textrm{ and }\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}>\frac{\sqrt{\log T}K^{1/3}}{T^{1/3}}\\ \end{cases}$
The next lemma states that irrespective of the state where you start from, if you alternate between two fixed arms for many rounds, then the state converges to a closed form solution that involves $\lambda$ and the $b_{i}$’s of the two alternating arms.

### v4

@@ B.3 Generalization for Unknown Replenishing Arm
Set $\varepsilon,\delta,M$ as stated in Theorem B.3.
Initialize rounds $t=1$.
/* Explore IV rewards and build their estimators: $\{\widehat{r}_{i}\}_{i\in[K]}$/
@@
2       Initialize reward estimate $\widehat{r}_{i}=0$.
3       for blocks $j\in[M]$ do // Restore the state to at least $b_{z}-\varepsilon$
             Choose an arm $z\in[K]$ uniformly at random.
              // $z$ = benchmark arm for state.
@@
5                   Play arm $z$.
6                   Update $t\leftarrow t+1$.
            Play arm $i$, observe reward $R_{j}^{i}$, and update: $\widehat{r}_{i}\leftarrow\widehat{r}_{i}+\frac{R_{j}^{i}}{M}$.
              // Play $i$ when $q\approx b_{z}-\varepsilon$.
@@
9       Initialize state estimator $\widehat{v}_{i}=0$.
10       for pulls $1,\dots,N(\lambda)$ do
11             Play arm $i$.
12             Update $t\leftarrow t+1$.
13
@@
             Play arm $i$, observe reward $S_{j}^{i}$, and update: $\widehat{v}_{i}\leftarrow\widehat{v}_{i}+\frac{S_{j}^{i}}{M}$.
              // Play $i$ when $q\approx b_{i}$
15
@@
17
Play arm $i_{R}$ for $N_{R}$ rounds, updating $t\leftarrow t+1$ after each one.
  // Restore state to at least $1-\varepsilon$
@@ Appendix C Appendix for Section 4
Input. $\eta=0.95\sqrt{\frac{\log K}{KT}}$, $\gamma=1.05\sqrt{\frac{K\log K}{T}},\beta=\sqrt{\frac{\log(K\delta^{-1})}{KT}},\forall\delta\in(0,1)$
Initially at time $t=1$, let $p_{1}$ be the uniform distribution over $[K]$.
for $t\in[T]$ do
       Choose an arm $I_{t}\in[K]$ from probability distribution $p_{t}$.
       Observe reward $g_{I_{t},t}\sim\textsf{Bern}(r_{I_{t}}q_{t})$.
       For each arm $i\in[K]$ compute the estimated biased gain:
@@
and update the estimated cumulative gain: $\tilde{G}_{i,t}=\sum_{s\in[t]}\tilde{g}_{i,s}.$
       Compute the new probability distribution over arms $p_{i,t+1}$:
@@ E.1 Noise-Perturbed Model
Note, that even in the noise-perturbed model the $q_{t}$ part of the state is still defined deterministically and Lemma 3.6 (and Corollary 3.7) still hold. Thus, after playing an arm $i$ repeatedly for $N(\lambda)$ rounds we get $|q_{t}-b_{i}|\leq\varepsilon$. Algorithm 2 plays replenishing arm $i_{R}$ for $N(\lambda)$ rounds and so $q_{t}\geq 1-\varepsilon$. After the state has converged to $b_{i_{R}}$, the algorithm starts building the estimators. While noise does not affect the $q_{t}$ part of the state, it does affect $\tilde{q}_{t}$ which in turn is where the estimators are built from. More specifically, for the $\hat{r}_{i}$ estimators, Algorithm 2 after playing $i_{R}$ for $N(\lambda)$ rounds, takes a sample of arm $i$; hence, the sample came from $\textsf{Bern}(r_{i}\cdot(1-\varepsilon+\nu_{t}))$. As a result, Lemmas 3.8 and 3.10 change as follows:
@@
The intuition and analysis of this case is similar to the analysis for the case where $\lambda=1-\epsilon$ (Section 5). If the noise added to a round $\nu_{t}$ is “small enough” (specifically, if $\nu_{t}\leq\sigma\sqrt{\log(T/\delta)}$ for $\sigma<\sqrt{T}$), then the analysis is identical to the case where $\lambda=1-\epsilon$. If the noise added to a round $\nu_{t}$ is greater than $\sigma\sqrt{\log(T/\delta)}$, then in the worst case the expected reward can be affected by a $\sigma T$ factor in total. Putting everything together (and using the Hoeffding bound of Equation 28) we get the result. ∎
@@ E.2 Unknown $\lambda$
In this section, we present the analysis for the case where $\lambda$ is unknown, i.e., the learner does not have any information about its value or the region that belongs to. When $\lambda$ is unknown, we can provide a general algorithm with sublinear regret (or sublinear approximate regret), when $\lambda\in[0,1]\setminus[\tilde{\Theta}(1/T),\omega((K/T)^{2/3})]$. The reason why our algorithm cannot recognize values of $\lambda$ in the interval $[\tilde{\Theta}(1/T),\omega((K/T)^{2/3})]$ (and hence, cannot provably achieve sublinear regret) will become clear after we introduce the algorithm. Formally, in this section we prove:
Theorem E.2.
Algorithm 8 for B-DES is agnostic to $\lambda$ and incurs regret:
  1. 1.
    If $\lambda>\omega((K/T)^{2/3})$: $R_{\textsc{DES}}\leq\mathcal{O}\left(\left(\frac{K\log(T)\log(\lambda)}{\log(1-\lambda)}\right)^{1/3}\cdot T^{2/3}\right)$.
  2. 2.
    If $\lambda\in[0,\tilde{\Theta}(1/T)]$:
    $R_{\textsc{DES}}(T)=\begin{cases}\widetilde{\mathcal{O}}\left(K^{1/3}T^{2/3}\right)&\textrm{for }\lambda\in[0,\Theta(1/T^{2})]\\ \mathcal{O}\left(T^{b/a}\right)&\textrm{for }\lambda=T^{-a/b},b<a<2b\\ (1-1/e)\textsf{OPT}&\textrm{for }\lambda=\tilde{\Theta}(1/T)\end{cases}$
The analysis is based on two algorithms: our main algorithm (B-DES Unknown $\lambda$, Algorithm 8) and an auxiliary algorithm called the $\lambda$-Estimator Builder Algorithm 7.
E.2.1 Algorithm Overview
At a high level, Algorithm 8 first identifies whether $\lambda$ is small (i.e., $\lambda\leq O((K/T)^{2/3})$ or large (i.e., $\lambda>\omega((K/T)^{2/3})$). Observing $\lambda$ directly is impossible. To overcome this, we pull a specific sequence of arms, study the structural properties of the rewards (and specifically, the difference of rewards) observed, and deduce $\lambda$; we refer to this process as obtaining “samples” of $\lambda$.
Let us be more specific: in the beginning, Algorithm 8 pulls each arm $i\in[K]$ for $(T/K)^{2/3}$ (rounds. Note that if $\lambda$ is large, then this means that we have just brought the state approximately at $b_{i}$ (Lemma 3.6). Once we have controlled the state, we draw $O((T/K)^{2/3})$ samples from $\textsf{Bern}(r_{i}q_{t})\approx\textsf{Bern}(r_{i}q_{i})$ and build the reward estimator $\hat{\mu}_{i}$; to do so, we draw $O((T/K)^{2/3})$ samples from $\textsf{Bern}(r_{i}q_{t})$, and compute the average reward.
Once we have estimators $\hat{\mu}_{i},\forall i\in[K]$, then we pick a random arm $j\in[K]$. For all arms $i\in[K]$ that are different than the randomly chosen arm $j$, we are going to play arms $i,j$ alternately for $O((T/K)^{2/3})$ rounds. Note that if $\lambda$ is big enough, after alternative between arms $i$ and $j$ for $O((T/K)^{2/3})$ rounds, the state has (roughly) converged to $b_{i}\frac{1-\lambda}{2-\lambda}+b_{j}\frac{1}{2-\lambda}$ (formally stated and proven in Lemma E.3). After controlling the state, the algorithm alternates between arms $i$ and $j$ for another $O((T/K)^{2/3})$ rounds, only logging the reward observed when arm $i$ is pulled (denoted by $R_{t}$ in the Algorithm): we call this estimator $\hat{\mu}_{i,j}$.
Why did we build both $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$? The difference between $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ helps us estimate whether $\lambda$ is “big” or “small”. Specifically, we distinguish two cases:
Case I.
If $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|\leq O(K^{1/3}T^{-1/3}\sqrt{\log T})$ holds for all arms, then:
  1. 1.
    Either $\lambda$ is very small (and so the state did not actually change enough in this interval),
  2. 2.
    or the $r_{i}$’s are very small $(\forall i\in[K])$,
  3. 3.
    or the values of the $b_{i}$’s are very close to each other.
More formally, that means that either $\max_{i,j\in[K]}\{r_{i}\cdot|b_{i}-b_{j}|\}\leq O(K^{1/3}T^{-1/3}\sqrt{\log T})$, or $\lambda\in[0,\tilde{\Theta}(1/T)]$. If the $r_{i}$’s are very small, then the regret of any algorithm that plays any random arm will be small as well. If the differences $|b_{i}-b_{j}|,\forall(i,j)\in[K]^{2}$ are very small, then that means that playing any arm will have similar effects to the value of the state. Hence, playing the best fixed arm (in terms of rewards $r_{i}$) is near optimal; in that case, playing EXP3.P will be able to yield sublinear regret. If we are in Case I (a), then EXP3.P is also the one that we use to obtain sublinear regret (see Section 4).
Case II.
If $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|$ is larger than $\omega(K^{1/3}T^{-1/3}\sqrt{\log T})$, then this means that $\lambda$ is “big” and that $\max_{i,j\in[K]}\{r_{i}\cdot|b_{i}-b_{j}|\}>\omega(K^{1/3}T^{-1/3}\sqrt{\log T})$. When we have identified that $\lambda$ is “big”, the next step is to construct an estimate for $N(\lambda)$, $\tilde{N}$. As we prove, if we can identify a $\tilde{N}$ such that $\tilde{N}=\Theta(N(\lambda))$ (specifically, we can estimate $\tilde{N}$ such that $N(\lambda)/\log T\leq\tilde{N}\leq 2N(\lambda)$), then we can essentially obtain an estimate $\hat{\lambda}$ of $\lambda$ such that: $|\hat{\lambda}-\lambda|\leq O(\varepsilon)$ with high probability (Lemma E.9). To obtain $\hat{\lambda}$, we use Algorithm 7.
At a high level, Algorithm 7 takes as input a pair of arms $i,j$ and the value of $\tilde{N}$ and creates estimators for quantities $r_{i}b_{i}$, $r_{i}b_{j}$ and $r_{i}b_{i}\frac{1-\lambda}{2-\lambda}+r_{i}b_{j}\frac{1}{2-\lambda}$; this is done in the same way that the main Algorithm 2 builds estimators. [^7] Recall that the only information that Algorithm 2 needs to know to build estimators is $N(\lambda)$. Note that $W_{i,j}(\lambda)\triangleq b_{i}\frac{1-\lambda}{2-\lambda}+b_{j}\frac{1}{2-\lambda}$ is the value that the state converges to after alternating between arms $i,j$ for a sufficient number of rounds. Let $\hat{r}_{i,i},\hat{r}_{i,j},$ and $\hat{r}_{i,j}^{\text{ALT}}$ be the estimators for quantities $r_{i}b_{i},r_{i},b_{j}$, and $r_{i}W_{i,j}(\lambda)$ respectively. Notice that we can solve a simple system of equations and obtain an estimate of $\lambda$ as:
$\hat{\lambda}=\frac{\hat{r}_{i,i}+\hat{r}_{i,j}-2\hat{r}_{i,j}^{\text{ALT}}}{\hat{r}_{i,i}-\hat{r}_{i,j}^{\text{ALT}}}.$
The only remaining piece is to explain how the estimate $\tilde{N}$ is obtained from Algorithm 8; our algorithm uses a doubling trick in order to estimate $\tilde{N}$. It starts with $\tilde{N}=1$. For each $\tilde{N}$, it first plays the anchoring arm $j$ for $\tilde{N}$ rounds, so that the state can be brought to be far enough from $b_{i}$. Then, it pulls arm $i$ for $\tilde{N}$ times, obtains a sample from $\textsf{Bern}(r_{i}\cdot q_{t})$, and repeats this process for $(T/K)^{2/3}$ rounds. If $\tilde{N}$ is big enough to move the value of the state near $b_{i}$, then the estimator $\tilde{\mu}_{i}$ should be near $r_{i}b_{i}$. As a result, the difference $|\hat{\mu}_{i}-\tilde{\mu}_{i}|$ will be small, and Algorithm 7 is called.
Now we present both algorithms:
Input. $\tilde{N}$, arm $i$, arm $j$.
Choose $\varepsilon$, $\delta$, $M$ as stated in Theorem 3.1.
Update $\tilde{N}\leftarrow\tilde{N}\log T$. // $\tilde{N}$ must be $\geq N(\lambda)$, but we only know that $\tilde{N}\geq N(\lambda)/\log T$
/* Build estimators for $r_{i}b_{j}$/
@@ E.2.1 Algorithm Overview
2       for pulls $\in[\tilde{N}]$ do
             Play arm $j$.
              // $q_{t}$ near $b_{j}$
@@
      Play arm $i$, observe reward $R_{i,j}$ and update: $\hat{r}_{i,j}=\hat{r}_{i,j}+R_{i,j}/M$.
        // $R_{i,j}\sim\textsf{Bern}(r_{i}b_{j})$
@@
5 for pulls $\in[\tilde{N}]$ do
       Play arm $i$.
        // $q_{t}$ near $b_{i}$
6
7for blocks $\in[M]$ do
       Play arm $i$, observe reward $R_{i,i}$ and update: $\hat{r}_{i,i}=\hat{r}_{i,i}+R_{i,i}/M$.
        // $R_{i,i}\sim\textsf{Bern}(r_{i}b_{i})$
8
/* Build estimators for $r_{i}W_{i,j}(\lambda)$/
9 for pulls $\in[\tilde{N}]$ do
@@ 10       Play arm $i$
11       Play arm $j$
12for blocks $\in[M]$ do
       Play arm $i$, observe reward $R_{i,j}^{\text{ALT}}$ and update: $\hat{r}_{i,j}^{\text{ALT}}=\hat{r}_{i,j}^{\text{ALT}}+R_{i,j}^{\text{ALT}}/M$.
        // $R_{i,j}^{\text{ALT}}\sim\textsf{Bern}(r_{i}W_{i,j}(\lambda))$
13       Play arm $j$.
14Solve $\hat{\lambda}=\frac{\hat{r}_{i,i}+\hat{r}_{i,j}-2\hat{r}_{i,j}^{\text{ALT}}}{\hat{r}_{i,i}-\hat{r}_{i,j}^{\text{ALT}}}$
Output. $\hat{\lambda}$.
ALGORITHM 7 $\lambda$-Estimator Builder
for each arm $i\in[K]$ do
       Play arm $i$ for $(T/K)^{2/3}$ rounds. /* If $\lambda$ is large: state now is at $q_{t}\approx b_{i}$. */
1       for $s=1,...,(T/K)^{2/3}$ rounds do
             Play arm $i$ and update estimator $\hat{\mu}_{i}\leftarrow\hat{r}_{i}+\frac{R_{s,i}}{(T/K)^{2/3}}$. // $R_{s,i}\sim\textsf{Bern}(r_{i}q_{t})$
2
3Choose a random arm $j\in[K]$.
4 for each $i\in[K]\setminus j$ do
5       Play arms $i,j$ alternately for $(T/K)^{2/3}$ rounds.
6       for $s=1,\dots,(T/K)^{2/3}$ rounds do
             Play arm $i$ and update $\hat{\mu}_{i,j}\leftarrow\hat{\mu}_{i,j}+\frac{R_{s,i}^{\text{ALT}}}{(T/K)^{2/3}}$. // $R_{s,i}^{\text{ALT}}\sim\textsf{Bern}(r_{i}q_{t})$
7             Play arm $j$.
8
9Choose $i\leftarrow\arg\max_{i\in[K]}|\hat{\mu}_{i}-\hat{\mu}_{i,j}|$.
/* Small $\lambda$. */
10 if $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|\leq 5\frac{\sqrt{\log T}}{(T/K)^{1/3}}$ then
11       Call EXP3.P
12
/* Large $\lambda$; ‘‘Binary search’’ for $N(\lambda)$/
13 else
14       $\tilde{N}\leftarrow 1$
15       $\tilde{\mu}_{i}\leftarrow\hat{\mu}_{i}+1$
16       while $|\hat{\mu}_{i}-\tilde{\mu}_{i}|>3\frac{\sqrt{\log T}}{(T/K)^{1/3}}$ do
17             Initialize estimator: $\tilde{\mu}_{i}\leftarrow 0$
18             for $(T/K)^{2/3}$ rounds do
                   Play arm $j$ for $\tilde{N}$ rounds. // Send state to a ‘‘common’’ baseline state
19                   Play arm $i$ for $\tilde{N}+1$ rounds.
                   Update $\tilde{\mu}_{i}\leftarrow\tilde{\mu}_{i}+\frac{R_{t}}{(T/K)^{2/3}}$ // $R_{t}\sim\textsf{Bern}(r_{i}q_{t})$; $q_{t}$ is similar $\forall$ samples (because of baseline from Line 12)
20
21            $\tilde{N}\leftarrow 2\cdot\tilde{N}$
22      Call Algorithm 7 for $\tilde{N}$, arms $i,j$ and get $\hat{\lambda}$.
23       Call Algorithm 2 for $\lambda=\hat{\lambda}$.
ALGORITHM 8 B-DES Unknown $\lambda$
E.2.2 Analysis
The next lemma states that irrespective of the state where you start from, if you alternate between two fixed arms for many rounds, then the state converges to a closed form solution that involves $\lambda$ and the $b_{i}$’s of the two alternating arms. For ease of analysis we define $W_{i,j}(\lambda)=b_{i}\frac{1-\lambda}{2-\lambda}+b_{j}\frac{1}{2-\lambda}$.